OpenStudy (anonymous):

what is an equation of the line with the slope 2 and passing through (2,7)?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

the possible answers given are y=2x-1, y=x+2, y=2x+3, y=2x+7, or none of t˙e answers are correcy

OpenStudy (anonymous):

can u plug in the numbers into the equation i gave u?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

no i cannot plug the numbers in. Its just a multiple choice question with only 5 possible answers

OpenStudy (anonymous):

lol y-y1=m(x-x1) you have to put the giving numbers into that equation to get the answer...if u try i will help u ; if u just need the answer im sorry

OpenStudy (anonymous):

oh okay i did that the first time lol i got 0=0

OpenStudy (anonymous):

show me how?

OpenStudy (anonymous):

7-7=2(2-2) 0=2(0) 0=0

OpenStudy (anonymous):

idk if im doing this right

OpenStudy (anonymous):

y-7=2(x-2) now distribute

OpenStudy (anonymous):

i just figured it out by distributing. Thank you so much

OpenStudy (anonymous):

which answer did u get

OpenStudy (anonymous):

i got y=2x+3